log5 (to the power of 5) (x+3)+log5(big 4)=log5(big 36)
once again its:
log5(x+3)+log5 4= log5 36

Answer by Alan3354(30993)

All logs are the same base, 5. The sum of logs on the left side means the 2 terms are multiplied. That is, the sum of 2 logs is the log of the product of the 2 terms.
So, log(x+3) + log(4) = log(36) (Regardless of base as long as they match)
log[(x+3)*4] = log(36)
If the logs are equal, the terms are equal.
(x+3)*4 = 36
4x+12 = 36
4x = 24
x = 6
Check: (6+3)*4 = 36.